OpenStudy (anonymous):

Use the quadratic formula to find the x-intercepts of the function. Round to the nearest tenth, if necessary. y = 2x2 + 7x + 5

OpenStudy (anonymous):

@timo86m

OpenStudy (luigi0210):

Do you know the quadratic formula?

OpenStudy (anonymous):

x intercepts is just another word for zero or roots make y=0 y = 2x2 + 7x + 5 0 = 2x2 + 7x + 5

OpenStudy (anonymous):

nope

OpenStudy (jdoe0001):

do you know what a y-intercept is?

OpenStudy (anonymous):

0 = (2)x2 + (7)x + (5) a b c 2 7 5 http://www.purplemath.com/modules/quadform.htm plug thos numbers in the formula

OpenStudy (anonymous):

X intercepts not y @jdoe0001

OpenStudy (jdoe0001):

hmm, pretty sure I read it correctly earlier on, yes, it does say NOW x-intercept

OpenStudy (anonymous):

http://www.purplemath.com/modules/quads/qform01.gif

OpenStudy (anonymous):

@jdoe0001 when I clicked on tim's link all it showed was a logo will you help me walk though it

OpenStudy (jdoe0001):

\(\bf \text{quadratic formula}\qquad \qquad x= \cfrac{ - b \pm \sqrt { b^2 -4ac}}{2a}\\ \begin{matrix} &a& b & c\\ y = &2x^2 &+ 7x& + 5 \end{matrix}\\ \qquad \qquad \qquad \qquad \qquad \qquad x= \cfrac{ - (7) \pm \sqrt { (7)^2 -4(2)(5)}}{2(2)}\)

OpenStudy (anonymous):

okay....

OpenStudy (jdoe0001):

so the root will give you 2 values for "x", those 2 are the solution or so-called zeros

OpenStudy (anonymous):

okay I think im still with you

OpenStudy (jdoe0001):

so, what did you get? :)

OpenStudy (anonymous):

just a sec and ill tell you:)

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

I got 27

OpenStudy (jdoe0001):

27?

OpenStudy (jdoe0001):

from \(x= \cfrac{ - (7) \pm \sqrt { (7)^2 -4(2)(5)}}{2(2)} \ \ ?\) hardly

OpenStudy (anonymous):

yea I put two in the places of x

OpenStudy (jdoe0001):

hmm actually you're meant to simplify only the right-hand side

OpenStudy (anonymous):

Oooo

OpenStudy (jdoe0001):

\notice the \(\huge \pm \)

OpenStudy (jdoe0001):

it means 2 solutions, one using the +, another using the -